### v1

# Gravitational collapse of a degenerate wormhole

Abstract
The dynamics of a degenerate spherically symmetric wormhole in a vacuum is considered. An extension of the equivalence principle to matter-free objects that are the source of a gravitational field is proposed. Using the Klinkhamer metric as an example, it is shown that a degenerate wormhole is precisely such an object. Application of the extended equivalence principle reduces the radial dynamics of the Klinkhamer wormhole to the dynamics of the radial fall of a test particle in a Schwarzschild gravitational field. It is proven that any bound state of the traversable Klinkhamer wormhole eventually collapses into a non-traversable Einstein-Rosen wormhole. An estimate is presented showing that the traversable Klinkhamer wormhole, although non-stationary, is a long-lived state.
keywords:
wormhole , degenerate metric , gravitational collapse
journal: Physics Letters B
The ideal gravitational collapse process is traditionally understood as the radial compression of self-gravitating dusty matter (Tolman, 1934; Oppenheimer and Snyder, 1939). In astrophysics, this corresponds to the final phase of black hole formation, when gravitational interaction becomes dominant. This picture is known to be approximate; a description of the real gravitational collapse process must take into account the pressure of matter (Shapiro and Teukolsky, 1983; Rezzolla and Zanotti, 2013; Özel and Freire, 2016). In this paper, we consider the gravitational collapse of a topological object free of matter – a degenerate wormhole. Unlike the gravitational collapse of matter, here the collapsing object is the wormhole throat, and the only interaction is indeed the gravitational interaction.
1 Introduction
In the pioneering work of Einstein and Rosen (Einstein and Rosen, 1935), it was shown that the source of the gravitational field, in addition to matter, is also curved space (i.e., the gravitational field itself). In this work, they constructed the first wormhole metric—the Einstein-Rosen bridge, connecting two sheets of the two-sheeted space. The Einstein-Rosen metric was degenerate, since its determinant $g$ vanishes on the spherical boundary that divides the two-sheeted space into two sheets.
This circumstance made it impossible to describe the degenerate metric within the framework of standard Einstein equations, since the determinant $g$ appears as a denominator in Christoffel symbols—and hence in the components of the Riemann curvature. Einstein and Rosen solved this problem by multiplying the standard Einstein equations by $g^{2}$. After this regularization, the Einstein equations become polynomial in the components of the metric tensor and their derivatives, allowing degenerate metrics to be included in the description. Subsequently, the regularized form of the Einstein equations was called polynomial and was formalized within the framework of the Einstein-Hilbert variational approach (Peres, 1963; Katanaev, 2006).
Since the degenerate Einstein-Rosen metric describes an empty and static curved space, the corresponding wormhole is an example of a matter-free stationary source of a gravitational field.
For a long time, interest in wormholes remained purely academic. However, the work of (Morris and Thorne, 1988) initiated the rapid development of concepts in this area (Visser, 1996). The main goal of this research was to study the conditions for the existence and stability of a traversable wormhole. The starting point for this direction was Thorne’s ”engineering” approach, which first takes a suitable wormhole metric and then determines the required matter distribution. A necessary element of this approach is the definiteness of the Riemann curvature, which is only possible with a non-degenerate metric. Therefore, non-degenerate wormhole states have long been (and remain) the primary object of study. The main obstacle in this direction was the inevitability of NEC violations (Hochberg and Visser, 1998), which requires the presence of exotic matter to stabilize a traversable wormhole. Note that the proof of this constraint is entirely based on the applicability of standard Einstein equations and, consequently, on the non-degeneracy of the wormhole metric.
Recent works by Klinkhamer (Klinkhamer, 2022, 2023, 2025; Wang, 2023) suggest a return to degenerate wormholes, where it all began. The Klinkhamer wormhole metric is a direct generalization of the Einstein-Rosen metric. The generalization is that the radius $b$ of a Klinkhamer wormhole can be larger than the gravitational radius, which ensures traversability. Constructed in complete analogy with the Einstein-Rosen metric, the Klinkhamer metric is also degenerate. Therefore, it should also be described not by the standard Einstein equations but by regularized equations—for the same reason that Einstein and Rosen did so in 1935. Like the Einstein-Rosen wormhole, the Klinkhamer wormhole is a matter-free source of a gravitational field, i.e. it possesses gravitational mass. By the equivalence principle, this must also imply that the Klinkhamer wormhole possesses an inertial mass.
This naturally raises the question: can such an object be stationary? The reason for this question is purely physical: as a ”self-gravitating” object, the Klinkhamer wormhole should exhibit a tendency to contract, i.e. to gravitational collapse. This allows us to consider the dynamics and gravitational collapse of a degenerate wormhole within a specific and well-controlled model. This problem is interesting both in itself and in light of a recent criticism of the Klinkhamer wormhole (Feng, 2023), which points to the lack of uniqueness of the solution to the Cauchy problem in this class of metrics.
Thus, the subject of this work is the dynamics of a degenerate wormhole using the Klinkhamer wormhole as an example.
The paper is organized as follows. In the second section, we consider the regularized Einstein equations and the Klinkhamer metric as one possible stationary solution of this system of equations. In the third section, we generalize the stationary Klinkhamer metric to a general non-stationary solution of the regularized equations, describing arbitrary wormhole dynamics within the framework of spherical symmetry. In the fourth section, we extend the equivalence principle to matter-free objects, an example of which is a degenerate wormhole. In the next, fifth section, we directly apply the extended equivalence principle, which allows us to reduce the dynamics of a spherically symmetric wormhole to the dynamics of a single test particle in a Schwarzschild gravitational field. In the sixth section, we present a well-known solution to the problem of the radial fall of a test particle in a Schwarzschild gravitational field. This allows us to obtain a unique solution to the problem of the dynamics of a degenerate spherical wormhole under given initial conditions in the seventh section, in particular, a complete solution to the gravitational collapse problem. The collapse results in the transformation of the Klinkhamer wormhole into an Einstein-Rosen wormhole. In the final, eighth part, we summarize and draw conclusions.
2 Stationary solution
We begin by presenting the Klinkhamer metric describing a degenerate spherical wormhole in a vacuum. We choose the Schwarzschild metric as the initial one [^1] The relativistic system of units is used, in which the speed of light $c$ and the gravitational constant $G$ are equal to 1, as well as the usual spherical coordinates $(r,\theta,\varphi)$, $d\Omega^{2}=sin^{2}\theta d\varphi^{2}+d\theta^{2}$ is an element of solid angle, $t$ is time, $M$ is the mass of the source of the gravitational field.
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\left(1-\frac{2M}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (1)
It is a solution to the standard Einstein equations in a vacuum [^2] The standard notation is adopted: $g_{\mu\nu}$ is the metric tensor, $R_{\mu\nu}$ and $R$ are the Ricci tensor and its trace.
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=0,$ (2)
defined in a one-sheeted space with a non-negative radial coordinate $r\geq 0$. We apply the topological dressing procedure of the two-sheeted coordinate transformation (Dimaschko, 2024) to the initial metric (1). To do this, we choose the following coordinate transformation
$r=\sqrt{l^{2}+b^{2}},$ (3)
which transforms the original one-sheeted space into the two-sheeted space of the wormhole. The new variable $l$ covers all values from -$\infty$ to $\infty$, with positive values $l>0$ corresponding to the first sheet, negative values $l<0$ to the second sheet of the two-sheeted space, and the value $l=0$ to the transition surface between them (i.e., the spherical throat $r=b$). The constant parameter $b\geq 2M$ represents the radius of the throat. The topological dressing transforms the original solution (1) of the standard Einstein equations (2) into a degenerate Klinkhamer metric (Klinkhamer, 2022)
$\begin{split}ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\left(1-\frac{2M}{r}\right)^{-1}\frac{l^{2}}{r^{2}}dl^{2}+r^{2}d\Omega^{2}.\end{split}$ (4)
This new metric, defined on a two-sheeted space, is a solution of the regularized Einstein equations in vacuum
$g^{2}\left(R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R\right)=0$ (5)
Despite the absence of matter, the degenerate wormhole (4) itself is the source of a gravitational field, which outside the throat is described by the Schwarzschild metric. To verify this, it is sufficient to convert the metric (4) to ordinary radial variables using the inverse of transformation (3):
$l=\pm\sqrt{r^{2}-b^{2}}\quad(r\geq b),$ (6)
Here, the choice of the “+” or “-” sign corresponds to the first or second sheet, respectively. After this transition to the usual $r$-coordinate, metric (4) returns to the form of the Schwarzschild metric (1), but with the constraint $r\geq b$:
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\left(1-\frac{2M}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2},\quad(r\geq b).$ (7)
This means that in the entire domain of definition $r\geq b$ of the metric (7), it coincides with the original Schwarzschild metric (1) – i.e., a matter-free throat creates on each of the two sheets the same gravitational field as a spherically symmetric distribution of matter with a total mass $M$.
3 Non-stationary solution
Thus, the Klinkhamer metric (4) describes a massive wormhole with mass $M$. This should lead to the instability of such a wormhole with respect to gravitational collapse, during which the radius of the throat $b$ decreases.
We will construct a non-stationary solution describing this process. To do this, in the original Schwarzschild metric (1), we perform a double coordinate substitution $(t,r)\mapsto(b,l)$ as follows:
$r=\sqrt{l^{2}+b^{2}},\quad t=t(b).$ (8)
Here $t(b)$ is a function that must be determined later. This substitution, in addition to the two-sheeted transformation of the radial coordinate as in (3), also transforms the radius $b$ of the wormhole into a new time variable. In the new variables, the non-stationary metric has the form
$ds^{2}=-\left(1-\frac{2M}{r}\right)\dot{t}^{2}\,db^{2}+\left(1-\frac{2M}{r}\right)^{-1}\frac{\left(l\,dl+b\,db\right)^{2}}{r^{2}}+r^{2}d\Omega^{2},$ (9)
where $\dot{t}=dt/db$. This metric describes a wormhole of variable radius, which is given by the function $b(t)$, the inverse of $t(b)$. Obtained from the Schwarzschild metric (1) by a coordinate transformation, this metric, like the stationary Klinkhamer metric (4), is a solution to the regularized Einstein equations (5). This follows directly from the covariance of Einstein’s equations and is surprisingly completely independent of the choice of the function $b(t)$.
This independence from the function $b(t)$ is because Einstein’s equations, which govern only the local metric, are not sufficient to fully describe the dynamics of space with a changing boundary [^3] The recognition that the wormhole throat constitutes an independent topological variable has precedent in the thin-shell wormhole literature (Visser, 1989, 1996), where the throat radius enters as the configuration variable in a Wheeler-De Witt quantization. The dynamical or statistical treatment of the topological variables has long been considered in discussions of topology change (Geroch, 1967; Sorkin, 1986; Horowitz, 1991; de Borde, 1994), wormhole formation (Hawking, 1990; Visser, 1996; Koga, 2023), and topological quantum field theory (Atiyah, 1989; Witten, 1988). More recently, (Borissova and Magueijo, 2025) derived an effective action for throat dynamics from first principles, showing that the throat radius is ’a single degree of freedom’ governed by its own equation of motion derived from the Israel junction formalism. (Dimaschko, 2024) extended this concept to degenerate wormholes, where the standard Einstein equations are insufficient. It was also shown that a direct application of the principle of least action allows one to include topological degrees of freedom in the analysis.. As shown in (Dimaschko, 2024), for such a complete description, the principle of least action must be directly used. On the one hand, this principle provides the Einstein’s equations themselves, which determine the local geometry of the space. On the other hand, this same principle determines the global topology of the space and, in particular, the dynamics of the wormhole boundary.
A similar approach could be taken to determine the dependence $b(t)$ that describes the collapse of a traversable wormhole. However, here we will do this more simply – directly using the equivalence principle.
4 Extended equivalence principle
When describing the dynamics of a degenerate wormhole under the influence of its own gravitational field, it is important to understand the source of this field. Consider this using the example of the Klinkhamer wormhole.
To do this, let us compare the Schwarzschild metric (1) and the Klinkhamer metric (7), written in the same $r$-coordinates. Both metrics contain the same mass parameter $M$, but its meaning in these two metrics is different. In the Schwarzschild metric, it is the mass of the centrally symmetric distribution of matter in 3D space. The Klinkhamer metric is a solution of the regularized Einstein equations (5) in the entire two-sheeted space, so matter is absent here. In this case, the source of the field is not matter, but the 2D surface of the two-sheeted space—its spherical throat $r=b$. Due to the symmetry of the two-sheet metric (4) with respect to the permutation of the first and second sheets (i.e., the change in the sign of the two-sheet coordinate $l$), this gravitational field is the same on both sheets.
Thus, the wormhole turns out to be an object endowed with a gravitational mass $M$. Does this object possess inertia? Without an answer to this question, it is impossible to construct the dynamics of a wormhole. To answer this question, we must turn to the equivalence principle, which postulates ”the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature of the body(Misner et al., 1973). Usually, the term ”body” implies that matter fills a region of 3D space. In the case of a wormhole, we are dealing with a different object—a compact gravitating 2D surface free of matter. Therefore, in this case, the equivalence principle must be expanded to ”the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature and dimensionality of the object.” This allows us to include not only matter but also objects that are self-consistent solutions of Einstein’s equations in a vacuum. We will call this formulation of the equivalence principle ”extended”.
How acceptable is such an extension? A necessary condition for its acceptability is the absence of internal contradictions and the ”integration” of its consequences into an already accepted and tested system of concepts and physically measurable quantities.
In considering the radial dynamics of the Klinkhamer wormhole, we will proceed from this extended interpretation of the equivalence principle. In any case, this cannot lead to contradictions with traditional general relativity, which is based on the standard equivalence principle. In studying the radial dynamics of the wormhole, we will demonstrate that this interpretation leads to natural results that are consistent with our understanding and are entirely analogous to the conventional picture of gravitational collapse.
5 Gravitational collapse and the extended equivalence principle
In this section, we show that a direct application of the extended equivalence principle reduces the field problem of the radial dynamics of a centrally symmetric wormhole (here, the Klinkhamer wormhole) to a single-particle problem of the motion of a test particle in a given gravitational field.
Note first that the use of new ($b,l)$ coordinates instead of the usual variables $(t,r)$ according to substitution (8) are nothing more than a basis for representing the non-stationary solution in the simple form (9). This representation is convenient for verifying the mathematical conformity of the solution with Einstein’s equations. On the other hand, for the physical analysis and interpretation of the solution, it is more convenient to use the usual variables $(t,r)$, which is what we will do next. In these ordinary coordinates, the wormhole surface appears simply as a sphere of variable radius $r=b(t)$, and the non-stationary metric (9) again has the form of a Schwarzschild metric with the constraint
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\left(1-\frac{2M}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2},\quad(r\geq b(t)).$ (10)
Let the radius of a collapsing wormhole decrease according to the law $b(t)$ from an initial value $b(0)>2M$. Consider the free motion of a test particle of mass $m$, located at the initial time $t=0$ near the wormhole surface, at a point with radial coordinate $r(0)=b(0)+0$. The test particle is assumed to be initially at rest relative to the surface of the wormhole.
By virtue of the extended equivalence principle, the position of the test particle relative to the wormhole surface should remain unchanged during the collapse process (the ”falling elevator effect”). This means that $b(t)$ must be identical to the law of radial motion of a particle $r(t)$ in a given gravitational field of the wormhole (see Fig. 1).
Thus, the problem of wormhole collapse is reduced to the problem of free fall of a test particle located near the surface of a collapsing wormhole.
Figure 1: By virtue of the extended equivalence principle, the collapse of a degenerate wormhole of mass $M$ is synchronous with the free fall of a test particle of mass $m$ located near the surface of the wormhole.
6 Free fall of a test particle
The free fall of a test particle of mass $m$ can be described by the Hamilton-Jacobi equation for a single-particle action $S$:
$g^{\mu\nu}\frac{\partial S}{\partial x^{\mu}}\frac{\partial S}{\partial x^{\nu}}=m^{2}.$ (11)
In the case of Schwarzschild geometry (1) and radial motion of the test particle of interest here, it has the form
$\left(1-\frac{2M}{r}\right)^{-1}\left(\frac{\partial S}{\partial t}\right)^{2}-\left(1-\frac{2M}{r}\right)\left(\frac{\partial S}{\partial r}\right)^{2}=m^{2}.$ (12)
We seek the solution of this equation in the form
$S=-\varepsilon t+S_{r}(r),$ (13)
where $\varepsilon$ is the total energy of the particle in the gravitational field. This yields an integral expression for $S_{r}(r)$:
$S_{r}(r)=\pm\int\left[\left(1-\frac{2M}{r}\right)^{-2}\varepsilon^{2}-\left(1-\frac{2M}{r}\right)^{-1}m^{2}\right]^{1/2}dr.$ (14)
Here, the choice of sign determines the direction of motion of the test particle relative to the radial axis. The condition $\partial S/\partial\varepsilon=0$ yields an integral dependence of $t$ on $r$
$t=\pm\frac{\varepsilon}{m}\int\left(1-\frac{2M}{r}\right)^{-1}\left[\left(\frac{\varepsilon}{m}\right)^{2}-\left(1-\frac{2M}{r}\right)\right]^{-1/2}dr,$ (15)
which can be transformed into a differential equation
$\frac{dr}{dt}=\pm\left(1-\frac{2M}{r}\right)\left[1-\left(\frac{m}{\varepsilon}\right)^{2}\left(1-\frac{2M}{r}\right)\right]^{1/2}.$ (16)
This equation determines the relationship between the radial velocity of a test particle and its radial coordinate. We further use this relationship to construct a phase portrait of a degenerate wormhole, describing, in particular, its gravitational collapse.
7 Radial dynamics of Klinkhamer’s wormhole
We have shown that, based on the extended equivalence principle, the dependence of the wormhole radius on time $b(t)$ should be identical to the dependence $r(t)$ for a test particle that falls freely. This means that the equation for the radius of the wormhole $b(t)$ is obtained from equation (16) by replacing $r(t)$ with $b(t)$:
$\frac{db}{dt}=\pm\left(1-\frac{2M}{b}\right)\left[1-\left(\frac{M}{E}\right)^{2}\left(1-\frac{2M}{b}\right)\right]^{1/2}.$ (17)
Here, the minus sign corresponds to the case of gravitational collapse (a decrease in the radius $b$), and the plus sign corresponds to the opposite case of inertial expansion (an increase in the radius $b$). Simultaneously, we replace the mass $m$ of the test particle with the mass $M$ of the wormhole and the total energy $\varepsilon$ of the test particle in a given gravitational field with the self-energy $E$ of the wormhole. The correctness of this substitution is verified by the Newtonian limit, when the collapse velocity is small $db/dt=v\ll 1$, and the wormhole radius $b$ is much larger than the gravitational radius $2M/b\ll 1$. Then, equation (17) leads to a non-relativistic expression for the total energy in a central gravitational field:
$E=M+\frac{Mv^{2}}{2}-\frac{M^{2}}{b}=\text{const},$ (18)
as expected [^4] A small but revealing subtlety: Newton’s expression (18) contains the potential energy of a spherical wormhole as $M^{2}/b$, while the Newtonian potential energy of a hollow sphere of radius $b$ is half as large. The reason for this is that for a hollow sphere in an ordinary one-sheeted space, this energy can be obtained by integrating the Newtonian energy density $\mathfrak{g}^{2}/(8\pi),\>\mathfrak{g}=M/r^{2}$ over the outer region $r>b$. In a two-sheeted space, the result of the same integration is twice as large, since it covers not one, but two sheets..
Equation (17) completely determines the phase portrait of the Klinkhamer wormhole. This phase portrait in coordinates $(b,db/dt)$ is shown below in Fig. 2. The separatrix here is the phase trajectory for which $E=M$. In the phase plane, it bounds the region within which $E<M$, and therefore the radius of the wormhole can vary only within finite limits (bound states).
Overall, the resulting phase portrait describes both the gravitational contraction of the wormhole (the lower part of the phase plane – below the separatrix) and its inertial expansion (the upper part of the phase plane – above the separatrix). In the first case, all phase trajectories converge at the point $(2M,0)$, corresponding to a non-traversable Einstein-Rosen wormhole. In the second case, the asymptotic limit of the process is the uniform inertial expansion of a traversable wormhole.
In the intermediate region, bounded by the separatrix, the radius of the wormhole varies within finite limits and culminates in a gravitational collapse at the point $(2M,0)$.
The process of gravitational collapse of the Klinkhamer wormhole into a non-traversable Einstein-Rosen wormhole, corresponding to trajectories in the lower half of the phase portrait, can be conveniently represented as the evolution of a Flamm paraboloid describing the wormhole. This evolution is shown in Fig. 3.
Figure 2: Phase portrait of the Klinkhamer wormhole dynamics. The lower part of the phase portrait describes the gravitational contraction of the wormhole, and the upper part describes its inertial expansion. The separatrix, shown by the thick line, bounds the region $E<M$, where the wormhole radius can vary only within finite limits. In this region, as in the region of gravitational contraction, the evolution of the Klinkhamer wormhole always ends with gravitational collapse into an Einstein-Rosen wormhole.
Figure 3: During gravitational collapse, the traversable Klinkhamer wormhole $(b>2M)$ transforms into the non-traversable Einstein-Rosen wormhole $(b=2M)$. The Flamm paraboloid has a kink at the throat, which disappears as a result of collapse (i.e., as $b\rightarrow 2M$). The three Flamm paraboloids shown are assumed to have identical values of the mass parameter $M=1$ and successively decreasing values of the throat radius: $b=6$, $b=4$, and $b=2$.
As shown in Appendix A, the Flamm paraboloid of the Klinkhamer metric has a kink at the throat. During gravitational collapse, the kink decreases with the throat radius until it completely disappears in the Einstein-Rosen wormhole limit (see Fig. 3).
The physical picture that emerges is intuitive: a traversable degenerate wormhole, lacking any matter to support it, is gravitationally unstable. Under its own gravitational attraction, the throat contracts until it reaches the minimum radius $b=2M$, at which point the wormhole becomes an Einstein-Rosen bridge—the ’maximally collapsed’ configuration of a degenerate wormhole. This final state is stable against further contraction since $b=2M$ corresponds to the Schwarzschild horizon, where the collapse velocity vanishes.
In conclusion of this section, let us dwell in more detail on the physical interpretation of the phase portrait shown in Fig. 2. This phase portrait describes three possible regimes:
1) collapse from an unbound state with $E>M$ (phase trajectories below the separatrix);
2) collapse from a bound state with $E<M$ (phase trajectories inside the separatrix);
3) the regime of unlimited expansion with $E>M$ (phase trajectories above the separatrix).
In each of these three regimes, the Klinkhamer wormhole, although non-stationary, nevertheless remains traversable for a certain time. This time depends on the initial conditions and can be quite significant. For example, the collapse from a resting state with an initial radius $b(0)\equiv b_{0}\gg 2M$ takes a time $\tau$ equal to the fall time of a test particle from a distance $b_{0}$ in the gravitational field of a classical point source of mass $M$. This time is (Landau and Lifshitz, 1982)
$\tau=\frac{\pi}{2}b_{0}\sqrt{\frac{b_{0}}{GM}}.$ (19)
For example, for the Klinkhamer wormhole with radius $b_{0}=10$ m and mass $M=10^{3}$ kg, the gravitational collapse time is $\tau=1.9\cdot 10^{5}$ s, i.e., approximately 2 days. Of course, the constructive value of this estimate is limited, as the question of the mechanism by which the initial state of a wormhole of radius $b_{0}$ is formed remains unanswered. However, this estimate shows that the Klinkhamer wormhole, although non-stationary, is a relatively long-lived object. It does not disappear ”at the speed of light.”
To study the conditions for the formation of a state with a degenerate metric, this metric must be linked to a specific physical process involving matter. For example, in astrophysics, this could involve the evolution of a neutron star, the loss of stability of a stellar state, the formation of a supernova, etc. To account for the possibility of states with a degenerate metric, these processes must be considered within the framework of regularized Einstein equations — but not the standard Einstein equations, as is currently the case.
8 Conclusions
Thus, using the Klinkhamer wormhole as an example, we have described the radial dynamics of a degenerate, spherically symmetric wormhole. A special case of this dynamic is gravitational collapse, transforming the Klinkhamer wormhole into a non-traversable Einstein-Rosen wormhole.
The result obtained here should be compared with Feng’s objection (Feng, 2023), who pointed out the apparent inconsistency of the non-stationary generalization of the Klinkhamer metric - and, consequently, of the Klinkhamer metric itself. Feng noted that the function $\lambda(t)$ (our $b(t)$) is not bounded by the Einstein equations, which makes the initial-value problem ill-posed and seemingly deprives the Klinkhamer metric of any physical meaning.
The resolution of the contradiction is the following:
1) the Einstein equations do indeed allow an arbitrary choice of the function $b(t)$. Feng is absolutely correct in this, and this is fully confirmed in Section 3 of our work;
2) however, not every choice of $b(t)$ describes the real dynamics of the wormhole.
The reason is that the local metric $g_{\mu\nu}$ alone does not provide a complete description of the wormhole. The complete description is given by the pair $(\mathscr{M},g_{\mu\nu})$, which, along with the local metric $g_{\mu\nu}$, additionally includes the global characteristics of the manifold $\mathscr{M}$. In (Dimaschko, 2024), it is shown that, with this in mind, applying the principle of least action simultaneously yields both Einstein’s equations for $g_{\mu\nu}$ and separate equations for the topological characteristics of the manifold $\mathscr{M}$. In the problem of the radial collapse of a wormhole, the only such characteristic is $b(t)$, the radius of the wormhole.
Due to spherical symmetry, the solution to this problem is significantly simplified by the possibility of applying the equivalence principle - extended to account for the presence of matter-free sources of the gravitational field. Direct application of the extended equivalence principle reduces the radial collapse problem of the Klinkhamer wormhole to the well-known problem of a test particle falling in a Schwarzschild gravitational field. This uniquely determines the dynamical equation for $b(t)$. As follows from the phase portrait in Fig. 2, the corresponding Cauchy problem has a unique solution for any initial conditions ($b_{0},\dot{b_{0}}$).
Thus, the contradiction discovered by Feng is resolved.
The correspondence established in this work between the radial dynamics of the Klinkhamer wormhole and the fall of a test particle in a Schwarzschild gravitational field allows us to determine the time of gravitational collapse. Simple estimates show that the Klinkhamer wormhole, although non-stationary, is a relatively long-lived state, realizing the idea of a traversable wormhole.
Appendix A Flamm’s paraboloid of the Klinkhamer metric
The general definition of Flamm paraboloid for an arbitrary spherically symmetric metric follows from the local Pythagorean relation, which expresses the square of the differential of the proper radial coordinate $dl^{2}$ through the sum of the squares of the differentials of the ordinary $r$-coordinate $dr^{2}$ and the auxiliary $z$-coordinate $dz^{2}$:
$dl^{2}=dr^{2}+dz^{2}$ (20)
Taking into account (9), the square of the differential of the proper radial coordinate is
$dl^{2}=\left(1-\frac{2M}{r}\right)^{-1}dr^{2},$ (21)
which yields the differential equation
$dz=\pm\left[\left(1-\frac{2M}{r}\right)^{-1}-1\right]^{1/2}\,dr.$ (22)
The solution, expressed as $r(z)$, determines the shape of the Flamm paraboloid for the Klinkhamer metric in cylindrical coordinates $(r,\varphi,z)$. For the unique determination of the solution $r(z)$, differential equation (A.3) requires a boundary condition, which we choose in the form
$r(0)=b.$ (23)
Its geometric meaning is that the throat, with radius b, corresponds to a cross-section of the paraboloid by the plane $z=0$. From this condition, we obtain
$r(z)=b+\frac{z^{2}}{8M}+\lvert z\rvert\sqrt{\frac{b}{2M}-1}$ (24)
The Flamm paraboloid of the Klinkhamer metric has a kink at the throat (that is, at $z=0$), which disappears at $b=2M$.
Declaration of Competing Interest
The author declares that he has no known competing financial interests or personal relationships that could influence the content of this paper.
Data availability
No data was used for the research described in the article.

### v2

@@
 Using the Klinkhamer metric as an example, it is shown that a degenerate wormhole is precisely such an object.
-Application of the extended equivalence principle reduces the radial dynamics of the Klinkhamer wormhole to the dynamics of the radial fall of a test particle in a Schwarzschild gravitational field.
+Applying the equivalence principle to matter-free field sources reduces the radial dynamics of the Klinkhamer wormhole to the dynamics of the radial fall of a test particle in a Schwarzschild gravitational field.
 It is proven that any bound state of the traversable Klinkhamer wormhole eventually collapses into a non-traversable Einstein-Rosen wormhole.
@@
 Since the degenerate Einstein-Rosen metric describes an empty and static curved space, the corresponding wormhole is an example of a matter-free stationary source of a gravitational field.
+In their original paper on the Einstein-Rosen bridge, its authors pointed out that the degenerate, geodesically complete solution they constructed of the regularized Einstein equations itself ”describes the field-producing mass without requiring the introduction of any new field quantities” (Einstein and Rosen, 1935), Section 1.
 For a long time, interest in wormholes remained purely academic.
@@
 Note that the proof of this constraint is entirely based on the applicability of standard Einstein equations and, consequently, on the non-degeneracy of the wormhole metric.
-Recent works by Klinkhamer (Klinkhamer, 2022, 2023, 2025; Wang, 2023) suggest a return to degenerate wormholes, where it all began.
+Recent works by Klinkhamer (Klinkhamer, 2023a, b, 2025; Wang, 2023) suggest a return to degenerate wormholes, where it all began.
 The Klinkhamer wormhole metric is a direct generalization of the Einstein-Rosen metric.
@@
 Constructed in complete analogy with the Einstein-Rosen metric, the Klinkhamer metric is also degenerate.
-Therefore, it should also be described not by the standard Einstein equations but by regularized equations—for the same reason that Einstein and Rosen did so in 1935.
-Like the Einstein-Rosen wormhole, the Klinkhamer wormhole is a matter-free source of a gravitational field, i.e.
-it possesses gravitational mass.
-By the equivalence principle, this must also imply that the Klinkhamer wormhole possesses an inertial mass.
+Therefore, it should also be described not by the standard Einstein equations but by regularized equations—for the same reason that Einstein and Rosen did so in 1935.Analogously to the Einstein–Rosen wormhole, the Klinkhamer wormhole constitutes a vacuum (purely geometric) solution that possesses an effective gravitational mass $M$, without the need to invoke any additional matter sources.
 This naturally raises the question: can such an object be stationary?
@@
 This problem is interesting both in itself and in light of a recent criticism of the Klinkhamer wormhole (Feng, 2023), which points to the lack of uniqueness of the solution to the Cauchy problem in this class of metrics.
-Thus, the subject of this work is the dynamics of a degenerate wormhole using the Klinkhamer wormhole as an example.
+This paper examines the dynamics of a degenerate wormhole using the example of the degenerate Klinkhamer wormhole. Specifically, the process of its gravitational collapse will be considered.
+Investigating the dynamical properties of the Klinkhamer wormhole necessitates an appropriate generalization of its original stationary metric to a broader class of nonstationary metrics, in such a way that the stationary configuration arises as a particular case.
+Describing the dynamics of a degenerate wormhole can be pursued through three possible approaches :
+1) by employing exclusively the regularized Einstein field equations themselves;
+2) by formulating and analyzing a reduced action, from which one derives the corresponding equations of motion for a suitably reduced set of dynamical variables;
+3) by constructing a reduced description grounded not on an action principle but rather on the equivalence principle, which likewise leads to a system of equations governing a reduced set of variables.
+Let us evaluate these three possibilities sequentially.
+1) Restricting the analysis solely to the regularized Einstein field equations is insufficient and reveals a fundamental limitation of this approach.
+The underlying reason is that the local metric $g_{\mu\nu}$ by itself does not furnish a complete characterization of the wormhole.
+A full description is provided instead by the pair $(\mathcal{M},g_{\mu\nu})$, which, in addition to the local metric $g_{\mu\nu}$, incorporates the global topological properties of the manifold $\mathcal{M}$.
+As demonstrated in (Dimaschko, 2024), once this fact is taken into account, the application of the principle of least action yields, in a unified manner, both the Einstein field equations for $g_{\mu\nu}$ and separate evolution equations for the topological characteristics of the manifold $\mathcal{M}$.
+It is precisely this circumstance that gives rise to the aforementioned Feng paradox.
+Specifically, in the context of the radial collapse of a degenerate wormhole, the sole topologically relevant degree of freedom is the time-dependent throat radius $b(t)$.
+The Einstein field equations are insensitive to the detailed temporal profile of $b(t)$; they impose only that the spacetime metric reduces to the Schwarzschild form in the exterior domain $r>b(t)$.
+As a result, the regularized Einstein equations, considered in isolation, are insufficient to capture the complete dynamical evolution of a degenerate wormhole.
+2) Employing a reduced action would be a natural strategy in this context; however, it results in a highly intricate mathematical problem.
+The underlying reason is that the Einstein–Hilbert action functional, based on the scalar curvature, becomes ill-defined when the metric is degenerate.
+Consequently, it is necessary to resort to a polynomial action (Peres, 1963; Katanaev, 2006), which possesses a substantially more involved structure.
+The formalism based on reduced variables is also well established within the thin-shell model, which arises as the zero-thickness limit of the Thorne framework as the layer thickness approaches zero Visser (1996).
+In this formulation, however, the spacetime metric remains nondegenerate and matter fields are present, thereby placing the Klinkhamer wormhole solution—characterized by a degenerate metric in the absence of matter—clearly outside scope of this model.
+3) Owing to its spherical symmetry, the degenerate Klinkhamer wormhole possesses a single dynamical degree of freedom, namely the wormhole radius $b$.
+Under these circumstances, the field dynamics effectively reduce to a one-dimensional problem, rendering it natural to invoke the equivalence principle.
+This principle is conventionally formulated in the context of point-particle dynamics and, to date, has not been systematically employed to construct the dynamics of field configurations.
+The Klinkhamer wormhole thus provides the first known example of a spacetime metric that permits the direct application of the equivalence principle in this manner, thereby mapping a field-theoretic problem onto an equivalent single-particle problem.
+This third approach is the one adopted in the present work.
+Its objective is to formulate the dynamical properties of a degenerate Klinkhamer wormhole on the basis of the equivalence principle.
+The proposed framework is predicated on extending the equivalence principle to field-theoretic (geometric) configurations that are free of matter sources.
+In the particular case of a degenerate wormhole, the central geometric object of interest is the wormhole throat.
 The paper is organized as follows.
@@
 In the fourth section, we extend the equivalence principle to matter-free objects, an example of which is a degenerate wormhole.
-In the next, fifth section, we directly apply the extended equivalence principle, which allows us to reduce the dynamics of a spherically symmetric wormhole to the dynamics of a single test particle in a Schwarzschild gravitational field.
+In the next, fifth section, we directly apply the equivalence principle, which allows us to reduce the dynamics of a spherically symmetric wormhole to the dynamics of a single test particle in a Schwarzschild gravitational field.
 In the sixth section, we present a well-known solution to the problem of the radial fall of a test particle in a Schwarzschild gravitational field.
-This allows us to obtain a unique solution to the problem of the dynamics of a degenerate spherical wormhole under given initial conditions in the seventh section, in particular, a complete solution to the gravitational collapse problem.
+In the seventh section, this allows us to obtain a unique solution to the problem of the dynamics of a degenerate spherical wormhole under given initial conditions, in particular, a complete solution to the gravitational collapse problem.
 The collapse results in the transformation of the Klinkhamer wormhole into an Einstein-Rosen wormhole.
+In the same section, we also examine the consistency between the equations of motion derived from the equivalence principle and the dynamical equations obtained in the Newtonian limit.
 In the final, eighth part, we summarize and draw conclusions.
@@
 We begin by presenting the Klinkhamer metric describing a degenerate spherical wormhole in a vacuum.
-We choose the Schwarzschild metric as the initial one [^1] The relativistic system of units is used, in which the speed of light $c$ and the gravitational constant $G$ are equal to 1, as well as the usual spherical coordinates $(r,\theta,\varphi)$, $d\Omega^{2}=sin^{2}\theta d\varphi^{2}+d\theta^{2}$ is an element of solid angle, $t$ is time, $M$ is the mass of the source of the gravitational field.
+We choose the Schwarzschild metric as the initial one [^1] The relativistic system of units is used, in which the speed of light $c$ and the gravitational constant $G$ are equal to 1, as well as the usual spherical coordinates $(r,\theta,\varphi)$, $d\Omega^{2}=d\theta^{2}+sin^{2}\theta d\varphi^{2}$ is an element of solid angle, $t$ is time, $M$ is the mass of the source of the gravitational field.
 $ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\left(1-\frac{2M}{r}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}.$ (1)
@@
 The constant parameter $b\geq 2M$ represents the radius of the throat.
-The topological dressing transforms the original solution (1) of the standard Einstein equations (2) into a degenerate Klinkhamer metric (Klinkhamer, 2022)
+The topological dressing transforms the original solution (1) of the standard Einstein equations (2) into a degenerate Klinkhamer metric (Klinkhamer, 2023a)
 $\begin{split}ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+\left(1-\frac{2M}{r}\right)^{-1}\frac{l^{2}}{r^{2}}dl^{2}+r^{2}d\Omega^{2}.\end{split}$ (4)
@@
 However, here we will do this more simply – directly using the equivalence principle.
-4 Extended equivalence principle
-When describing the dynamics of a degenerate wormhole under the influence of its own gravitational field, it is important to understand the source of this field.
-Consider this using the example of the Klinkhamer wormhole.
-To do this, let us compare the Schwarzschild metric (1) and the Klinkhamer metric (7), written in the same $r$-coordinates.
-Both metrics contain the same mass parameter $M$, but its meaning in these two metrics is different.
-In the Schwarzschild metric, it is the mass of the centrally symmetric distribution of matter in 3D space.
-The Klinkhamer metric is a solution of the regularized Einstein equations (5) in the entire two-sheeted space, so matter is absent here.
-In this case, the source of the field is not matter, but the 2D surface of the two-sheeted space—its spherical throat $r=b$.
-Due to the symmetry of the two-sheet metric (4) with respect to the permutation of the first and second sheets (i.e., the change in the sign of the two-sheet coordinate $l$), this gravitational field is the same on both sheets.
-Thus, the wormhole turns out to be an object endowed with a gravitational mass $M$.
-Does this object possess inertia?
-Without an answer to this question, it is impossible to construct the dynamics of a wormhole.
-To answer this question, we must turn to the equivalence principle, which postulates ”the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature of the body(Misner et al., 1973).
-Usually, the term ”body” implies that matter fills a region of 3D space.
-In the case of a wormhole, we are dealing with a different object—a compact gravitating 2D surface free of matter.
-Therefore, in this case, the equivalence principle must be expanded to ”the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature and dimensionality of the object.” This allows us to include not only matter but also objects that are self-consistent solutions of Einstein’s equations in a vacuum.
-We will call this formulation of the equivalence principle ”extended”.
-How acceptable is such an extension?
-A necessary condition for its acceptability is the absence of internal contradictions and the ”integration” of its consequences into an already accepted and tested system of concepts and physically measurable quantities.
-In considering the radial dynamics of the Klinkhamer wormhole, we will proceed from this extended interpretation of the equivalence principle.
-In any case, this cannot lead to contradictions with traditional general relativity, which is based on the standard equivalence principle.
-In studying the radial dynamics of the wormhole, we will demonstrate that this interpretation leads to natural results that are consistent with our understanding and are entirely analogous to the conventional picture of gravitational collapse.
+4 Extended interpretation of the equivalence principle
+In the analysis of the dynamics of a degenerate wormhole subjected to its self-generated gravitational field, it is essential to identify and characterize the source of this field.
+This issue can be elucidated concretely by considering the specific example of the Klinkhamer wormhole.
+To proceed, we compare the Schwarzschild metric (1) with the Klinkhamer metric (7), expressed in the same radial coordinate $r$.
+Both metrics involve the same mass parameter $M$, but its physical interpretation differs in the two cases.
+In the Schwarzschild metric, $M$ represents the mass of a spherically symmetric distribution of matter in three-dimensional space.
+By contrast, the Klinkhamer metric is an exact solution of the regularized Einstein equations (5) on the entire two-sheeted manifold, where no matter is present.
+In this latter case, the source of the gravitational field is not matter in the usual sense, but rather the two-dimensional surface associated with the two-sheeted geometry—specifically, its spherical throat at $r=b$.
+Owing to the symmetry of the two-sheet metric (4) under the exchange of the two sheets (equivalently, the change of sign of the two-sheet coordinate $l$), the resulting gravitational field is identical on both sheets.
+Thus, the wormhole can be regarded as an object endowed with a gravitational mass $M$.
+A natural question then arises: does this object also possess inertia?
+Without addressing this issue, it is not possible to formulate a consistent dynamical description of a wormhole.
+To resolve this, we invoke the equivalence principle, which postulates
+“the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature of the body(Misner et al., 1973).
+In conventional usage, the term “body” refers to a material distribution occupying a three-dimensional spatial region.
+In the wormhole case, however, the situation is fundamentally different: we are dealing with a compact, gravitating two-dimensional surface that is devoid of matter.
+Accordingly, the standard formulation of the equivalence principle must be generalized to assert
+the equality (in suitable units) of the inertial and gravitational masses, regardless of the nature and dimensionality of the object.
+What criteria determine the admissibility of such an extension?
+A necessary condition for its acceptability is the absence of internal inconsistencies, as well as the successful incorporation of its implications into an already established and empirically tested system of theoretical concepts and operationally defined, physically measurable quantities.
+In Section 7 of this paper, we show, that in the Newtonian limit, the extended formulation of the equivalence principle precisely corresponds to the principles of ordinary classical mechanics extended to a two-sheeted space.
+In this sense, it passes the test of the correspondence principle and allows for a natural extension to general relativity.
+Therefore, in our further consideration of the radial dynamics of the Klinkhamer wormhole, we will always proceed from this extended interpretation of the equivalence principle.
+Thus, we regard the extended equivalence principle as an axiom, on the same footing as the standard equivalence principle in general relativity, and will next test it by consistency and correspondence arguments.
 5 Gravitational collapse and the extended equivalence principle
-In this section, we show that a direct application of the extended equivalence principle reduces the field problem of the radial dynamics of a centrally symmetric wormhole (here, the Klinkhamer wormhole) to a single-particle problem of the motion of a test particle in a given gravitational field.
+In this section, we show that a direct application of the equivalence principle reduces the field problem of the radial dynamics of a spherically symmetric wormhole (here, the Klinkhamer wormhole) to a single-particle problem of the motion of a test particle in a given gravitational field.
 Note first that the use of new ($b,l)$ coordinates instead of the usual variables $(t,r)$ according to substitution (8) are nothing more than a basis for representing the non-stationary solution in the simple form (9).
@@
 The test particle is assumed to be initially at rest relative to the surface of the wormhole.
-By virtue of the extended equivalence principle, the position of the test particle relative to the wormhole surface should remain unchanged during the collapse process (the ”falling elevator effect”).
+By virtue of the equivalence principle, the position of the test particle relative to the wormhole surface should remain unchanged during the collapse process (the ”falling elevator effect”).
 This means that $b(t)$ must be identical to the law of radial motion of a particle $r(t)$ in a given gravitational field of the wormhole (see Fig.
@@
 Thus, the problem of wormhole collapse is reduced to the problem of free fall of a test particle located near the surface of a collapsing wormhole.
-Figure 1: By virtue of the extended equivalence principle, the collapse of a degenerate wormhole of mass $M$ is synchronous with the free fall of a test particle of mass $m$ located near the surface of the wormhole.
+[^4] In this case, the possibility of applying the equivalence principle is due to the absence of non-gravitational forces.
+In (Dimaschko, 2024), a charged wormhole was considered against the background of an electrovacuum.
+In this case, the equivalence principle was inapplicable due to the presence of non-gravitational forces.
+Therefore, to determine the equilibrium radius of the wormhole, a direct application of the principle of least action was required.
+Figure 1: By virtue of the equivalence principle, the collapse of a degenerate wormhole of mass $M$ is synchronous with the free fall of a test particle of mass $m$ located near the surface of the wormhole.
 6 Free fall of a test particle
@@
 We further use this relationship to construct a phase portrait of a degenerate wormhole, describing, in particular, its gravitational collapse.
-7 Radial dynamics of Klinkhamer’s wormhole
-We have shown that, based on the extended equivalence principle, the dependence of the wormhole radius on time $b(t)$ should be identical to the dependence $r(t)$ for a test particle that falls freely.
+7 Radial dynamics and self-energy of the wormhole
+We have shown that, based on the equivalence principle, the dependence of the wormhole radius on time $b(t)$ should be identical to the dependence $r(t)$ for a test particle that falls freely.
 This means that the equation for the radius of the wormhole $b(t)$ is obtained from equation (16) by replacing $r(t)$ with $b(t)$:
-$\frac{db}{dt}=\pm\left(1-\frac{2M}{b}\right)\left[1-\left(\frac{M}{E}\right)^{2}\left(1-\frac{2M}{b}\right)\right]^{1/2}.$ (17)
+$\frac{db}{dt}=\pm\left(1-\frac{2M}{b}\right)\left[1-\left(\frac{M}{\mathcal{E}}\right)^{2}\left(1-\frac{2M}{b}\right)\right]^{1/2}.$ (17)
 Here, the minus sign corresponds to the case of gravitational collapse (a decrease in the radius $b$), and the plus sign corresponds to the opposite case of inertial expansion (an increase in the radius $b$).
-Simultaneously, we replace the mass $m$ of the test particle with the mass $M$ of the wormhole and the total energy $\varepsilon$ of the test particle in a given gravitational field with the self-energy $E$ of the wormhole.
-The correctness of this substitution is verified by the Newtonian limit, when the collapse velocity is small $db/dt=v\ll 1$, and the wormhole radius $b$ is much larger than the gravitational radius $2M/b\ll 1$.
-Then, equation (17) leads to a non-relativistic expression for the total energy in a central gravitational field:
-$E=M+\frac{Mv^{2}}{2}-\frac{M^{2}}{b}=\text{const},$ (18)
-as expected [^4] A small but revealing subtlety: Newton’s expression (18) contains the potential energy of a spherical wormhole as $M^{2}/b$, while the Newtonian potential energy of a hollow sphere of radius $b$ is half as large.
-The reason for this is that for a hollow sphere in an ordinary one-sheeted space, this energy can be obtained by integrating the Newtonian energy density $\mathfrak{g}^{2}/(8\pi),\>\mathfrak{g}=M/r^{2}$ over the outer region $r>b$.
-In a two-sheeted space, the result of the same integration is twice as large, since it covers not one, but two sheets..
+Simultaneously, we replace the mass $m$ of the test particle with the mass $M$ of the wormhole and the total energy $\varepsilon$ of the test particle in a given gravitational field with the ”self-energy” $\mathcal{E}$ of the wormhole.
+In fact, the quantity $\mathcal{E}/M$, which appears in the new differential equation (17), inherits from the ratio $\varepsilon/m$, which enters the original differential equation (16), only its kinematic meaning—the meaning of the first integral of motion.
+This meaning, in complete analogy with (16), consists of nothing more than the conservation of a certain combination of the radius $b$ of the wormhole and the rate $db/dt$ of change of this radius:
+$\left(1-\frac{2M}{b}\right)^{-1}\left[1-\left(1-\frac{2M}{b}\right)^{-1}\left(\frac{db}{dt}\right)^{2}\right]=\text{const}=\left(\frac{M}{\mathcal{E}}\right)^{2}.$ (18)
+To clarify the independent dynamic meaning of the ”self-energy” $\mathcal{E}$, we will proceed in the same way as is customary to clarify the meaning of the formal parameter $M$ in the Schwarzschild solution: namely, we will move to the Newtonian limit.
+In the context of the equation of motion (18), this implies a lower approximation in the small parameter $2M/b$, within which equation (18) takes the form
+$\mathcal{E}=M+\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-\frac{M^{2}}{b}.$ (19)
+Let us compare this expression with the result that follows directly from Newton’s theory of the gravitational field for the same spherical wormhole of mass $M$ and radius $b$.
+This requires calculating the corresponding energy $U$ of the gravitational field, which in the Newtonian approximation is a locally well-determined quantity:
+$U=-\frac{1}{8\pi}\int\mathfrak{g}^{2}\,dV.$ (20)
+Here, $\mathfrak{g}$ is the local acceleration of gravity, which in the spherically symmetric case has an absolute value
+$g=\frac{M}{b^{2}},\quad(r\geq b),$ (21)
+and the integration extends to the entire two-sheeted space.
+This corresponds to a double integration of expression (20) over the region $r\geq b$ of the ordinary one-sheeted space and leads to the following expression for the total gravitational energy:
+$U(b)=-\frac{2}{8\pi}\int_{r\geq b}g^{2}\,dV=-\frac{M^{2}}{b}.$ (22)
+The factor 2 in front of the integral in this expression appears precisely because of the two-sheeted nature of space and the double integration over the region $r\geq b$ .
+After this, the Newtonian wormhole Lagrangian, written in the usual form,
+$L\!\left(b,\frac{db}{dt}\right)=\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-U(b),$ (23)
+immediately leads to the conservation of the total energy of the wormhole, which includes the energy of the field (22):
+$\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-\frac{M^{2}}{b}=\text{const.}$ (24)
+By denoting the constant on the right-hand side as EME-M, we arrive at the law of conservation of the total energy $E$ of the wormhole:
+$E=M+\frac{M}{2}\left(\frac{db}{dt}\right)^{2}-\frac{M^{2}}{b}.$ (25)
+This expression for the total energy $E$ coincides with the previous expression (19) for the formal parameter $\mathcal{E}$.
+Thus, in the Newtonian limit, the formal conserved quantity $\mathcal{E}$ arising in the equation of motion (18) naturally reproduces the wormhole’s self-energy $E$, which includes the field energy.
+This provides grounds for maintaining the same interpretation in the relativistic case, where the gravitational field energy is no longer a locally, but rather a globally defined quantity.
+For this reason, we shall henceforth refer to the corresponding integral of motion $\mathcal{E}$ as the wormhole’s self-energy, without any further qualifications.
+We have thus demonstrated that, in the Newtonian limit, the extended formulation of the equivalence principle we adopted is completely consistent with ordinary classical mechanics—both lead to identical equations of motion (19) and (25).
+It is precisely the correspondence principle that reduces the applicability of this extended formulation to the classical mechanics, which is initially based on our everyday experience.
+A rigorous derivation of the extended equivalence principle from the polynomial action is a much more difficult task and is left for future research.
+Another important observation following from this analysis is the significant influence of the two-sheeted space on the wormhole dynamics.
+The factor 2 in front of the integral in expression (22) is a direct consequence of the two-sheeted space of the wormhole.
+In a one-sheeted space, this factor would be equal to 1, which would lead to a half-expression for the field energy $U(b)$.
+This would cause a discrepancy between the equation of motion for a test particle, determined by the potential M/b-M/b, and the analogous equation of motion for the surface of a matter-free object in a one-sheeted space, determined by the half-potential M/(2b)-M/(2b).
+In turn, such a discrepancy would result in desynchronization of the motion of the test particle and the gravitating surface (see Fig.
+1), and to a violation of the equivalence principle in its extended formulation.
+This means that the extended formulation of the equivalence principle not only allows for a two-sheeted structure of space, but also requires it - as a necessary condition for the existence of matter-free sources of the gravitational field.
+In this sense, the extended formulation of the equivalence principle and the two-sheetedness of the wormhole space manifest themselves as mutually conditioned factors, which imparts rigidity to their combination in the theory of a degenerate wormhole.
+Of course, the example of a degenerate wormhole considered here has a limited degree of generality.
+However, it alone already indicates that, at least in the case of spherical symmetry,
+consistency of the equivalence principle with Newtonian mechanics requires a two-sheeted structure.
 Equation (17) completely determines the phase portrait of the Klinkhamer wormhole.
@@
 In the second case, the asymptotic limit of the process is the uniform inertial expansion of a traversable wormhole.
-In the intermediate region, bounded by the separatrix, the radius of the wormhole varies within finite limits and culminates in a gravitational collapse at the point $(2M,0)$.
-The process of gravitational collapse of the Klinkhamer wormhole into a non-traversable Einstein-Rosen wormhole, corresponding to trajectories in the lower half of the phase portrait, can be conveniently represented as the evolution of a Flamm paraboloid describing the wormhole.
-This evolution is shown in Fig.
-3.
 Figure 2: Phase portrait of the Klinkhamer wormhole dynamics.
@@
 In this region, as in the region of gravitational contraction, the evolution of the Klinkhamer wormhole always ends with gravitational collapse into an Einstein-Rosen wormhole.
+In the intermediate region, bounded by the separatrix, the radius of the wormhole varies within finite limits and culminates in a gravitational collapse at the point $(2M,0)$.
+Note that here the limit $b\rightarrow 2M$ does not produce the usual Schwarzschild interior region $0<r<2M$, but rather the Einstein–Rosen bridge connecting two copies of the exterior Schwarzschild region $r\geq 2M$ via the degenerate throat at $r=2M$.
+In the Einstein-Rosen-Horowitz-Klinkhamer polynomial formulation, this configuration is geodesically complete and free of curvature singularities on both sheets; the singular region $r=0$ of the standard maximally extended Schwarzschild solution is simply absent from the manifold.
+The process of gravitational collapse of the Klinkhamer wormhole into a non-traversable Einstein-Rosen wormhole, corresponding to trajectories in the lower half of the phase portrait, can be conveniently represented as the evolution of a Flamm paraboloid describing the wormhole.
+This evolution is shown in Fig.
+3.
 Figure 3: During gravitational collapse, the traversable Klinkhamer wormhole $(b>2M)$ transforms into the non-traversable Einstein-Rosen wormhole $(b=2M)$.
@@
 This final state is stable against further contraction since $b=2M$ corresponds to the Schwarzschild horizon, where the collapse velocity vanishes.
+By ‘stable final state’ we mean that the collapsing degenerate wormhole asymptotically approaches this static Einstein–Rosen configuration and cannot evolve further.
+No further contraction beyond $b=2M$ is possible because the manifold by construction has no region with $r<2M$.
 In conclusion of this section, let us dwell in more detail on the physical interpretation of the phase portrait shown in Fig. 2. This phase portrait describes three possible regimes:
@@
 For example, the collapse from a resting state with an initial radius $b(0)\equiv b_{0}\gg 2M$ takes a time $\tau$ equal to the fall time of a test particle from a distance $b_{0}$ in the gravitational field of a classical point source of mass $M$.
-This time is (Landau and Lifshitz, 1982)
+This time is (Landau and Lifshitz, 1982), Eq.(103.17):
 $\tau=\frac{\pi}{2}b_{0}\sqrt{\frac{b_{0}}{GM}}.$ (26)
@@
 Due to spherical symmetry, the solution to this problem is significantly simplified by the possibility of applying the equivalence principle - extended to account for the presence of matter-free sources of the gravitational field.
-Direct application of the extended equivalence principle reduces the radial collapse problem of the Klinkhamer wormhole to the well-known problem of a test particle falling in a Schwarzschild gravitational field.
+Direct application of the equivalence principle reduces the radial collapse problem of the Klinkhamer wormhole to the well-known problem of a test particle falling in a Schwarzschild gravitational field.
 This uniquely determines the dynamical equation for $b(t)$.